# Bi-directional Model Cascading with Proxy Confidence

### v1

@@ Appendix H Delegated Prompt Length
Appendix I Varying Forward Confidence Training Size
Table  10 shows the deferral AUC using Bi-Dir deferral while limiting the training data available to the auxiliary model. Deferral performance declines slightly under training data limitations in all deferral rate spans.
Table 10. Bi-directional deferral performance using smaller training sets on MMLU.
AUC (.2) AUC (.4) AUC (full)
Random 0.094603 0.194561 0.528991
Bi-Directional (Aux 25%) 0.097455 0.203643 0.552574
Bi-Directional (Aux 50%) 0.097741 0.204144 0.552880
Bi-Directional (full) 0.097874 0.204550 0.553650
Bi-Directional (Oracle Aux.) 0.099757 0.209051 0.558934
Appendix J Bootstrap Confidence Intervals
@@ Appendix I Varying Forward Confidence Training Size
Table 11. Bi-Dir AUC improvement over baseline (MaxProb) 95% bootstrap CI .
@@
Table 12. BackInt AUC improvement over baseline (MaxProb) 95% bootstrap CI .
@@
Appendix K Limitations
In this work we evaluate LLM cascades using only models from the LLaMA-2 series due to their recency, transparency and accessibility. While this choice reduces confounds and facilitates reproducibility, there are some limitations as a result. Larger models may better evaluate the value of cascades in very difficult tasks, the performance of our small model on difficult tasks shows a significant amount of “guessing”, while this is expected for the smaller model it introduces an element of chance that is not conducive to consistent results. There is also some evidence that diversity of component models in a cascade may improve its performance which we are unable to explore with this model set (Lebovitz et al., 2023).

### v2

@@ Appendix H Deferred Prompt Length
Appendix I Bootstrap Confidence Intervals
@@
Table 11: Bi-Dir AUC improvement over baseline (MaxProb) 95% bootstrap CI .
@@
Table 12: BackInt AUC improvement over baseline (MaxProb) 95% bootstrap CI .
@@
Appendix J Limitations
In this work we evaluate LLM cascades using only models from the LLaMA-2 series due to their recency, transparency and accessibility. While this choice reduces confounds and facilitates reproducibility, there are some limitations as a result. Larger models may better evaluate the value of cascades in very difficult tasks, the performance of our small model on difficult tasks shows a significant amount of “guessing”, while this is expected for the smaller model it introduces an element of chance that is not conducive to consistent results. There is also some evidence that diversity of component models in a cascade may improve its performance which we are unable to explore with this model set [24].